Multiprocessor Implementation of Nondeterminate Computation in a

Functional Programming Framework

Computation Structures Group Memo 375 January 5, 1995

Jack B. Dennis and Guang R. Gao

The research reported in this docment was performed, in part, using facilities of the Laboratory for Computer Science of the Massachusetts Institute of Technology. Funding for the Laboratory is provided in part by the Advanced Research Projects Agency of the Department of Defense under the Office of Naval Research contract N00014-92-J-1310.